OpenStudy (blahblahblah):

PLEASE HELP ME I'LL MEDAL AND FOLLOW A LUCK PERSON AND FOLLOW EVERYONE ELSE! Translate the following phrase to an expression. Then, evaluate the expression for n = 18. the product of a number and 2 18 n ; 36 2 n ; 36 n + 2; 20 n ÷ 2; 9

OpenStudy (nickbeengone):

2 n ; 36

OpenStudy (nickbeengone):

2*18=36 36;36 36=36

563blackghost (563blackghost):

`Product` means the multiplying of two numbers so it would be... \(\Large\bf{n \times 2}\) You would then input 18 as n since it is given `n = 18` \(\Large\bf{18 \times 2}\)

OpenStudy (nickbeengone):

Basically what I just said.